Rank the following pressures from of highest to lowest pressure.

1.45 atm

800 kPa

10.5 in Hg 1.20 bar
950 mmHg
Question 1 options:

A)
10.5 in Hg > 1.20 bar > 950 mmHg > 1.45 atm > 800 kPa

B)
950 mm Hg > 800 kPa > 10.5 in Hg > 1.45 atm > 1.20 bar

C)
800 kPa > 1.45 atm > 950 mm Hg > 1.20 bar > 10.5 in Hg

D)
1.45 atm > 10.5 in Hg > 950 mm Hg > 800 kPa > 1.20 bar

E)
1.20 bar > 800 kPa > 10.5 in Hg > 950 mm Hg > 1.45 atm

I want to the levels of pressure like which is the highest?
Answer #1
C)
800 kPa > 1.45 atm > 950 mm Hg > 1.20 bar > 10.5 in Hg

conversion 1atm = 101.325KPa = 760mmHg=1.01325bar
Answer #2

B)
950 mm Hg > 800 kPa > 10.5 in Hg > 1.45 atm > 1.20 bar
Answer #3

1.45 atm = 1.45*10^5 N/m^2

800 kPa = 800*10^3 N/m^2 =8*10^5 N/m^2

10.5 inch Hg = 10.5*25.4 mm Hg =266.7 mm Hg = (266.7/760) *10^5 N/m^2 =0.351*10^5 N/m^2

1.20 bar = 1.20 atm = 1.20*10^5 N/m^2

950 mm Hg = (950/760)*10^5 N/m^2 =1.25*10^5 N/m^2


Correct answer is 800kPa > 1.45 atm > 950 mmHg >1.20 bar >10.5 inch Hg